OpenStudy (anonymous):

A random variable, X, is such that X~ Bin(50,0.6). The random variable Y is the mean of 50 random measurements of X. What will be the distribution of Y?

OpenStudy (anonymous):

Y will have normal distribution

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

not really ><

OpenStudy (anonymous):

Mean=np=50*0.6=30 Variance=npq=50*0.6*(1-0.6)=12 Standard Deviation= Square root (12) Thus,